# Vitali Telezkia and Stefan Klumppa

### v1

@@ 3.4 Columnar clusters
At intermediate and high density and strong fields, the behavior of the system in dominated by oriented chains and bands. Snapshots of the system show additional structure in the scenarios classified as bands or oriented chains: figure 9 shows that as the magnetic field strength increases, a disordered network of chains first organizes into a broad band, which then, as the field strength is increased further, splits into parallel clusters of oriented chains. These structures are aligned in the direction of the external field and, for sufficiently strong magnetic interactions, span the whole system. This effect seems to be more pronounced at the intermediate density $\Phi=0.23$, on which we will focus in the following.
@@
Fig.  9: Snapshots of system with intermediate density and strong magnetic interactions show columnar clusters, parallel bands spanning the whole system in the direction of the field that shrink with increasing external field. The density is $\Phi=0.23$, the velocity $v=23$.
@@
Snapshots of the simulations indicate (see figure 9) that the number of lanes, i.e., of parallel chains in each cluster can vary and depends on the field strength. The clusters appear to be quite regularly spaced in the direction perpendicular to the external magnetic field. All these properties are reminiscent of magnetic field-induced columnar clusters in ferrofluids 22, 23, 24.
@@
Fig.  10: Left: snapshot of system at $B=14$, $\mu=2.6$, $v_{0}=14$ showing columnar clusters. Individual clusters are color coded, center of geometry is marked by a red cross. Right: average cluster size $\overline{N}$ against external magnetic field strength $B$ for particles with moderate self-propulsion speed $v_{0}=14$. System is at intermediate density $\Phi=0.23$. Colors indicate different magnetic moments of the dipolar particles. Data sampled over trajectories at times $t=18-20$ every $\Delta t=0.1$.
@@
We characterized these columnar clusters with two parameters: the average number of lanes in a cluster $\overline{n_{\mathrm{l}}}$ and the columnar spacing $\overline{\Delta r_{y}}$, which is the average distance between neighboring cluster centers perpendicular to the external magnetic field. These two parameters were calculated as follows for simulations showing oriented chains or band (i.e., for simulations with $\left\langle\phi_{e}\right\rangle>0.5$ and $\left\langle\phi_{p}\right\rangle>0.5$): First, we identified individual clusters (color coded in figure 10, left), using a density based cluster algorithm (see Methods). We then calculated the center of geometry (red crosses) and the length of each cluster in direction of the external magnetic field $l_{x}^{i}$. Averaging the distances between neighboring centers of geometry perpendicular to the direction of the external magnetic field while accounting for periodic boundary conditions, gives us the columnar spacing $\overline{\Delta r_{\mathrm{y}}}$. To approximate the number of lanes in each cluster, we divided the number of particles per cluster $n_{i}$ by the cluster extension in direction of the external magnetic field $n_{\mathrm{l}}^{i}=\frac{n_{i}}{l_{x}^{i}}$.
@@
Indeed, the average number of particles per cluster $\overline{N}$ (figure 10, right) decreases rapidly with increasing magnetic field strength, confirming that clusters separate into smaller sized clusters as the magnetic field strength increases in systems with $\mu>1$. The average cluster size appears to converge to a fixed value for very strong external magnetic fields.
@@
Fig.  11: Columnar spacing $\overline{\Delta r_{y}}$ (left) and average number of lanes per cluster $\overline{n_{\mathrm{l}}}$ (right) against external magnetic field strength $B$. The magnetic moment was set to $\mu=4$ and density to $\Phi=0.23$. Colors indicate increasing self-propulsion speed of the particles. Data sampled over trajectories at times $t=18-20$ every $\Delta t=0.1$. Data points show systems that fulfill criteria $\left\langle\phi_{e}\right\rangle>0.5$ and $\left\langle\phi_{p}\right\rangle>0.5$.
@@
Figure 11 shows the columnar spacing $\overline{\Delta r_{y}}$ and the average number of lanes per column $\overline{n_{\mathrm{l}}}$ for systems with strong magnetic interactions $\mu=4$. Both the columnar spacing and the number of lanes per cluster decay rapidly and approach a fixed value as the external magnetic field strength increases. Activity seems to have only a minor effect on this behavior. Columnar spacing has been reported to be independent of the external magnetic field in experiments with (passive) ferrofluids 24. Our result for active particles at for strong magnetic fields mirror this behavior. However, we observed a strong dependence of the columnar spacing in the regime of weak external magnetic fields. We note that in some cases, a weak external magnetic field seems to promote separation between columnar structures.
@@
As reported in 24, columnar spacing is reduced with increasing density in ferrofluids. While we did qualitatively observe in movies of our simulations that the number of lanes in each column increases with density, more systematic investigation of density effects would be needed to make quantitative statements about the influence of density on columnar structures.
4 Conclusions
@@
In this study, we used extensive simulations to investigat the effect of a homogeneous external magnetic field on systems of active dipolar particles. Based on the combination of three order parameters, we determied diagrams of state for different magnetic field strengths to provide a comprehensive overview of the collective behaviors observed in such system. Consistent with previous work 14, 15, we see that dipolar interactions without an external field result in chains, which at higher densities form networks and bands. Motility-induces phase separation is only observed in the absence of dipolar interactions, already the weakest interaction we simulated resulted in networks of chains instead. The dipolar interaction likely prevents the jammed configurations and, thus, the slow-down required for motility-induced phase separation.
@@
Under the influence of an external magnetic field oriented chains are dominant and bands are seen less frequently. For weak interactions, an oriented gas of individual dipoles is observed. Both states are only weakly dependent on the the level of activity and practically independent of it when a strong external magnetic field is present. This observation indicates that for strong field, the magnetic interactions dominate over activity, such that systems of active dipolar particles resemble passive ferrofluids more than classical (non-dipolar) active Brownian particles. This observation for strong field stands in stark contrast to the chain formation observed in systems without external magnetic fields, where the activity plays a crucial role in collective phenomena 14, 15, 18.
@@
Two observations may merit further research in future work: (i) We observed oscillating chains as a transient state towards band formation in our system. We suspect that these oscillations are caused by a buckling instability and that fragmentation of long chains due to buckling allows chains of different lengths to coexist in dilute systems, despite high magnetic moments. The self-assembled active dipolar chains we study here have some similarities to active polymers, in which the monomers are linked by a spring potential 43. In that case, buckling has been studied both theoretically and experimentally 44, 45, 46, 47 and was shown to be controlled by the flexure number, which characterizes the relation between active propulsion and bending rigidity 44, 45.
@@
(ii) We observed that active dipolar particles form columnar clusters that resemble structures induced by external magnetic fields in (passive) ferrofluids 22, 23, 24. A quantitative analysis showed that the number of lanes per column and the columnar spacing decreases rapidly with increasing magnetic field strength, but become constant for strong fields. The strong-field behavior is similar to observations in ferrofluids, where Columnar spacing has been reported to be independent of the external magnetic field strength 22, 24. Further features of the system and experimental protocols (such as the application of the external magnetic field, presence of hydrodynamic interactions, polydispersity and boundary conditions) may have an effect on this result. Here, in a first step, we have shown here activity does not strongly affect separation between columns. In future studies, the effect of density could be investigated in more detail and compared to observations for ferrofluids 22.
@@
Our results together with those from previous studies 14, 17 show that a preferred orientation (introduced by the external field) coupled to activity can significantly change the observed collective behaviors. In our model, activity is introduced as an active force acting in the direction of the magnetic moment of the particle. Therefore, orientation competes with the formation of structures based on random or conflicting directions of self-propulsion such as MIPS. These observations indicate that the details of how activity is introduced into the system and whether or how the direction of self-propulsion is coupled to the direction of the dipoles should be important for the type of patterns and collective phenomena that are observed. They suggest that even richer collective phenomena are possible if there is loose coupling between the dipolar and the self-propulsion direction of the active particles.
Conflicts of interest
There are no conflicts to declare.
Acknowledgements
@@
Notes and references
  • Vicsek and Zafeiris 2012 T. Vicsek and A. Zafeiris, Phys. Rep., 2012, 517, 71–140.
  • Romanczuk et al. 2012 P. Romanczuk, M. Bär, W. Ebeling, B. Lindner and L. Schimansky-Geier, Eur. Phys. J. Spec. Top., 2012, 202, 1–162.
  • Bechinger et al. 2016 C. Bechinger, R. Di Leonardo, H. Löwen, C. Reichhardt, G. Volpe and G. Volpe, Rev. Mod. Phys., 2016, 88, 45006.
  • Cates and Tailleur 2015 M. E. Cates and J. Tailleur, Annu. Rev. Condens. Matter Phys., 2015, 6, 219–244.
  • Bente et al. 2018 K. Bente, A. Codutti, F. Bachmann and D. Faivre, Small, 2018, 14, 1704374.
  • Klumpp et al. 2019 S. Klumpp, C. T. Lefèvre, M. Bennet and D. Faivre, Phys. Rep., 2019, 789, 1–54.
  • Bazylinski and Frankel 2004 D. A. Bazylinski and R. B. Frankel, Nat. Rev. Microbiol., 2004, 2, 217–230.
  • Baraban et al. 2012 L. Baraban, D. Makarov, R. Streubel, I. Monch, D. Grimm, S. Sanchez and O. G. Schmidt, ACS nano, 2012, 6, 3383–3389.
  • Alsaadawi et al. 2021 Y. Alsaadawi, A. Eichler-Volf, M. Heigl, P. Zahn, M. Albrecht and A. Erbe, The European Physical Journal E, 2021, 44, 23.
  • Ghosh and Fischer 2009 A. Ghosh and P. Fischer, Nano letters, 2009, 9, 2243–2245.
  • Vach et al. 2015 P. J. Vach, P. Fratzl, S. Klumpp and D. Faivre, Nano letters, 2015, 15, 7064–7070.
  • Carlsen et al. 2014 R. W. Carlsen, M. R. Edwards, J. Zhuang, C. Pacoret and M. Sitti, Lab on a Chip, 2014, 14, 3850–3859.
  • Magdanz et al. 2013 V. Magdanz, S. Sanchez and O. G. Schmidt, Advanced materials, 2013, 25, 6581–6588.
  • Liao et al. 2020 G. J. Liao, C. K. Hall and S. H. Klapp, Soft Matter, 2020, 16, 2208–2223.
  • Telezki and Klumpp 2020 V. Telezki and S. Klumpp, Soft Matter, 2020, 16, 10537–10547.
  • Obreque et al. 2024 P. M. Obreque, O. Garrido, D. Romero, H. Löwen and F. Guzmán-Lastra, arXiv preprint arXiv:2403.02569, 2024.
  • Sesé-Sansa et al. 2022 E. Sesé-Sansa, G.-J. Liao, D. Levis, I. Pagonabarraga and S. H. Klapp, Soft Matter, 2022, 18, 5388–5401.
  • Maloney et al. 2020 R. C. Maloney, G. J. Liao, S. H. Klapp and C. K. Hall, Soft Matter, 2020, 16, 3779–3791.
  • Maloney and Hall 2020 R. C. Maloney and C. K. Hall, Langmuir, 2020, 36, 6378–6387.
  • Waisbord et al. 2016 N. Waisbord, C. T. Lefèvre, L. Bocquet, C. Ybert and C. Cottin-Bizonne, Physical Review Fluids, 2016, 1, 053203.
  • Meng et al. 2018 F. Meng, D. Matsunaga and R. Golestanian, Physical review letters, 2018, 120, 188101.
  • Liu et al. 1995 J. Liu, E. M. Lawrence, A. Wu, M. L. Ivey, G. A. Flores, K. Javier, J. Bibette and J. Richard, Phys. Rev. Lett., 1995, 74, 2828.
  • Flores et al. 1999 G. A. Flores, J. Liu, M. Mohebi and N. Jamasbi, Phys. Rev. E - Stat. Physics, Plasmas, Fluids, Relat. Interdiscip. Top., 1999, 59, 751–762.
  • Ivey et al. 2001 M. Ivey, J. Liu, Y. Zhu and S. Cutillas, Phys. Rev. E - Stat. Nonlinear, Soft Matter Phys., 2001, 63, 011403.
  • Yung et al. 1998 K. W. Yung, P. B. Landecker and D. D. Villani, Magn. Electr. Sep., 1998, 9, 39–52.
  • Chandler et al. 1983 D. Chandler, J. D. Weeks and H. C. Andersen, Science (80-. )., 1983, 220, 787–794.
  • Yung et al. 1999 K. W. Yung, P. B. Landecker and D. D. Villani, Magn. Electr. Sep., 1999, 10, 29–33.
  • Anderson et al. 2008 J. A. Anderson, C. D. Lorenz and A. Travesset, J. Comput. Phys., 2008, 227, 5342–5359.
  • Anderson et al. 2020 J. A. Anderson, J. Glaser and S. C. Glotzer, Comput. Mater. Sci., 2020, 173, 109363.
  • Adorf et al. 2018 C. S. Adorf, P. M. Dodd, V. Ramasubramani and S. C. Glotzer, Comput. Mater. Sci., 2018, 146, 220–229.
  • Ramasubramani et al. 2018 V. Ramasubramani, C. Adorf, P. Dodd, B. Dice and S. Glotzer, Proc. 17th Python Sci. Conf., 2018, pp. 152–159.
  • Klapp 2005 S. H. L. Klapp, J. Phys. Condens. Matter, 2005, 17, R525—-R550.
  • Holm and Weis 2005 C. Holm and J.-J. J. Weis, Curr. Opin. Colloid Interface Sci., 2005, 10, 133–140.
  • Barupal and Fiehn 2019 D. K. Barupal and O. Fiehn, Environ. Health Perspect., 2019, 127, 2825–2830.
  • Bruss and Glotzer 2017 I. R. Bruss and S. C. Glotzer, Soft Matter, 2017, 13, 5117–5121.
  • Melle and Martin 2003 S. Melle and J. E. Martin, J. Chem. Phys., 2003, 118, 9875–9881.
  • Melle et al. 2003 S. Melle, O. G. Calderón, M. A. Rubio and G. G. Fuller, Phys. Rev. E - Stat. Physics, Plasmas, Fluids, Relat. Interdiscip. Top., 2003, 68, 041503.
  • De Vicente et al. 2011 J. De Vicente, D. J. Klingenberg and R. Hidalgo-Alvarez, Soft Matter, 2011, 7, 3701–3710.
  • Mohebi et al. 1996 M. Mohebi, N. Jamasbi and J. Liu, Phys. Rev. E - Stat. Physics, Plasmas, Fluids, Relat. Interdiscip. Top., 1996, 54, 5407–5413.
  • Weis 2003 J. J. Weis, J. Phys. Condens. Matter, 2003, 15, S1471.
  • Weis and Levesque 2005 J. J. Weis and D. Levesque, Adv. Polym. Sci., 2005, 185, 163–225.
  • Chaté et al. 2008 H. Chaté, F. Ginelli, G. Grégoire and F. Raynaud, Phys. Rev. E - Stat. Nonlinear, Soft Matter Phys., 2008, 77, 046113.
  • Winkler and Gompper 2020 R. G. Winkler and G. Gompper, The journal of chemical physics, 2020, 153, 040901.
  • Prathyusha et al. 2018 K. R. Prathyusha, S. Henkes and R. Sknepnek, Phys. Rev. E, 2018, 97, 022606.
  • Abbaspour et al. 2023 L. Abbaspour, A. Malek, S. Karpitschka and S. Klumpp, Physical Review Research, 2023, 5, 013171.
  • Kurjahn et al. 2023 M. Kurjahn, A. Deka, A. Girot, L. Abbaspour, S. Klumpp, M. Lorenz, O. Baumchen and S. Karpitschka, eLife, 2023, 12, RP87450.
  • Vliegenthart et al. 2020 G. A. Vliegenthart, A. Ravichandran, M. Ripoll, T. Auth and G. Gompper, Science advances, 2020, 6, eaaw9975.

### v2

@@ III.4 Columnar clusters
At intermediate and high density and strong fields, the behavior of the system in dominated by oriented chains and bands. Snapshots of the system show additional structure in the scenarios classified as bands or oriented chains: fig. 10 shows that as the magnetic field strength increases, a disordered network of chains first organizes into a broad band, which then, as the field strength is increased further, splits into parallel clusters of oriented chains. These structures are aligned in the direction of the external field and, for sufficiently strong magnetic interactions, span the whole system. This effect seems to be more pronounced at the intermediate density $\Phi=0.23$, on which we will focus in the following.
@@
Figure 10: Snapshots of system with intermediate density and strong magnetic interactions show columnar clusters, parallel bands spanning the whole system in the direction of the field that shrink with increasing external field. The density is $\Phi=0.23$, the velocity $v=23$.
@@
Snapshots of the simulations indicate (see fig. 10) that the number of lanes, i.e., of parallel chains in each cluster can vary and depends on the field strength. The clusters appear to be quite regularly spaced in the direction perpendicular to the external magnetic field. All these properties are reminiscent of magnetic field-induced columnar clusters in ferrofluids [22, 23, 24].
@@
Figure 11: Left: snapshot of system at $B=14$, $\mu=2.6$, $v_{0}=14$ showing columnar clusters. Individual clusters are color coded, center of geometry is marked by a red cross. Right: average cluster size $\overline{N}$ against external magnetic field strength $B$ for particles with moderate self-propulsion speed $v_{0}=14$. System is at intermediate density $\Phi=0.23$. Colors indicate different magnetic moments of the dipolar particles.
@@
We characterized these columnar clusters with two parameters: the average number of lanes in a cluster $\overline{n_{\mathrm{l}}}$ and the columnar spacing $\overline{\Delta r_{y}}$, which is the average distance between neighboring cluster centers perpendicular to the external magnetic field. These two parameters were calculated as follows for simulations showing oriented chains or band (i.e., for simulations with $\left\langle\phi_{e}\right\rangle>0.5$ and $\left\langle\phi_{p}\right\rangle>0.5$): First, we identified individual clusters (color coded in fig. 11, left), using a density based cluster algorithm (see Methods). We then calculated the center of geometry (red crosses) and the length of each cluster in direction of the external magnetic field $l_{x}^{i}$. Averaging the distances between neighboring centers of geometry perpendicular to the direction of the external magnetic field while accounting for periodic boundary conditions, gives us the columnar spacing $\overline{\Delta r_{\mathrm{y}}}$. To approximate the number of lanes in each cluster, we divided the number of particles per cluster $n_{i}$ by the cluster extension in direction of the external magnetic field $n_{\mathrm{l}}^{i}=\frac{n_{i}}{l_{x}^{i}}$.
@@
Indeed, the average number of particles per cluster $\overline{N}$ (fig. 11, right) decreases rapidly with increasing magnetic field strength, confirming that clusters separate into smaller sized clusters as the magnetic field strength increases in systems with $\mu>1$. The average cluster size appears to converge to a fixed value for very strong external magnetic fields.
@@
Figure 12: Columnar spacing $\overline{\Delta r_{y}}$ (left) and average number of lanes per cluster $\overline{n_{\mathrm{l}}}$ (right) against external magnetic field strength $B$. The magnetic moment was set to $\mu=4$ and density to $\Phi=0.23$. Colors indicate increasing self-propulsion speed of the particles. Data sampled over trajectories at times $t=18-20$ every $\Delta t=0.1$. Data points show systems that fulfill criteria $\left\langle\phi_{e}\right\rangle>0.5$ and $\left\langle\phi_{p}\right\rangle>0.5$.
@@
Figure 12 shows the columnar spacing $\overline{\Delta r_{y}}$ and the average number of lanes per column $\overline{n_{\mathrm{l}}}$ for systems with strong magnetic interactions $\mu=4$. Both the columnar spacing and the number of lanes per cluster decay rapidly and approach a fixed value as the external magnetic field strength increases. Activity seems to have only a minor effect on this behavior. Columnar spacing has been reported to be independent of the external magnetic field in experiments with (passive) ferrofluids [24]. Our result for active particles at for strong magnetic fields mirror this behavior. However, we observed a strong dependence of the columnar spacing in the regime of weak external magnetic fields. We note that in some cases, a weak external magnetic field seems to promote separation between columnar structures.
@@
As reported in [24], columnar spacing is reduced with increasing density in ferrofluids. While we did qualitatively observe in movies of our simulations that the number of lanes in each column increases with density, more systematic investigation of density effects would be needed to make quantitative statements about the influence of density on columnar structures.
IV Conclusions
@@
In this study, we used extensive simulations to investigate the effect of a homogeneous external magnetic field on systems of active dipolar particles. Based on the combination of three order parameters, we determined diagrams of state for different magnetic field strengths to provide a comprehensive overview of the collective behaviors observed in such system. Consistent with previous work [14, 15], we see that dipolar interactions without an external field result in chains, which at higher densities form networks and bands. Motility-induces phase separation is only observed in the absence of dipolar interactions, already the weakest interaction we simulated resulted in networks of chains instead. The dipolar interaction likely prevents the jammed configurations and, thus, the slow-down required for motility-induced phase separation.
@@
Under the influence of an external magnetic field oriented chains are dominant and bands are seen less frequently. For weak interactions, an oriented gas of individual dipoles is observed. Both states are only weakly dependent on the the level of activity and practically independent of it when a strong external magnetic field is present. This observation indicates that for strong field, the magnetic interactions dominate over activity, such that systems of active dipolar particles resemble passive ferrofluids more than classical (non-dipolar) active Brownian particles. This observation for strong field stands in stark contrast to the chain formation observed in systems without external magnetic fields, where the activity plays a crucial role in collective phenomena [14, 15, 18].
@@
Two observations may merit further research in future work: (i) We observed oscillating chains as a transient state towards band formation in our system. We suspect that these oscillations are caused by a buckling instability and that fragmentation of long chains due to buckling allows chains of different lengths to coexist in dilute systems, despite high magnetic moments. The self-assembled active dipolar chains we study here have some similarities to active polymers, in which the monomers are linked by a spring potential [43]. In that case, buckling has been studied both theoretically and experimentally [44, 45, 46, 47] and was shown to be controlled by the flexure number, which characterizes the relation between active propulsion and bending rigidity [44, 45].
@@
(ii) We observed that active dipolar particles form columnar clusters that resemble structures induced by external magnetic fields in (passive) ferrofluids [22, 23, 24]. A quantitative analysis showed that the number of lanes per column and the columnar spacing decreases rapidly with increasing magnetic field strength, but become constant for strong fields. The strong-field behavior is similar to observations in ferrofluids, where Columnar spacing has been reported to be independent of the external magnetic field strength [22, 24]. Further features of the system and experimental protocols (such as the application of the external magnetic field, presence of hydrodynamic interactions, polydispersity and boundary conditions) may have an effect on this result. Here, in a first step, we have shown here activity does not strongly affect separation between columns. In future studies, the effect of density could be investigated in more detail and compared to observations for ferrofluids [22].
@@
Our results together with those from previous studies [14, 17] show that a preferred orientation (introduced by the external field) coupled to activity can significantly change the observed collective behaviors. In our model, activity is introduced as an active force acting in the direction of the magnetic moment of the particle. Therefore, orientation competes with the formation of structures based on random or conflicting directions of self-propulsion such as MIPS. These observations indicate that the details of how activity is introduced into the system and whether or how the direction of self-propulsion is coupled to the direction of the dipoles should be important for the type of patterns and collective phenomena that are observed. They suggest that even richer collective phenomena are possible if there is loose coupling between the dipolar and the self-propulsion direction of the active particles.
Acknowledgements.